Journalist Report – April 17th

Crew 192 Journalist Report 17Apr2018

Author: Richard Blakeman

This outstanding crew of scientists has amazed me with their ability to constantly come up with new science experiments and projects in the wake of our original projects being lost.

Even when the weather is prohibitive for EVAs, they come up with ideas, experiments, educational discussions, and training sessions that we can perform in the confines of the hab, things like knot-tying lessons and other survival skills, tips on the use of social media in science outreach, insight into the psychology of astronaut group dynamics and crew selection, and more.

Very proud to be part of such a bright, dynamic and creative minds and I’m excited to continue working with them in the future.

Greenhab Report – April 17th

GreenHab Report
Name: Andreea Radulescu
Date: 17 Apr 2018

Environmental control:

Ambient with door open few hours
Shade cloth: on
Fan: not on.

Average temperatures: 21.1℃
Low 13.7℃
High 53.2℃

Hours of supplemental light: None

Changes to crops: None

Daily water usage for crops: 9 gal

Time(s) of watering for crops: Twice: In the morning around 9:30 am and afternoon 5:30pm.

Moringa research observations: None

Changes to research plants: No

Aquaponics: None

Narrative: No significant changes. Picked another batch of tomatoes for the crew as well as peppers and few baby carrots.

Support/supplies needed: None

Operations Report – April 17th

Crew 192 Operations Report 17Apr2018

SOL 9

Name of person filing report: Victoria Varone

Non-nominal systems: None

Generator (hours run): 12 hours

Solar ‒ SOC 87%

Diesel – 75% fuel

Propane – 62% (telemetry reported by Shannon) psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – All ATVs filled 100%, 25% left in one can, requesting gas refill.

Water (trailer) – Empty, tank being cleaned in RAM.
Water (static) – 730 gallons (two external tanks)
Trailer to Static Pump used – No
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 133616.7

Toilet tank emptied: No. Emptied yesterday, April 16.

ATV’s Used: Red ATV, Blue ATV’s #1, #2, and #3

Oil Added? No but checked and OK.

ATV Fuel Used: 20%

# Hours the ATVs were Used today: 6

Notes on ATVs: All ATVs working well, oil checked, gas filled today.

Deimos rover used: No
Hours: 120
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: No
Hours: 31.2 (corrected)
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Opportunity rover used: No
Hours: 29.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Curiosity rover used: No
Hours: 25.9
Beginning charge: 91%
Ending charge: 91%
Currently charging: Yes

HabCar used and why, where? No

General notes and comments: Crew is doing well, trying to make the best use of time for EVAs in the midst of constant strong winds.

Summary of internet: 0MB at 1900.

Summary of suits and radios: Suits all charging, three two-piece suits used, radios working fine. Suit #4 charging port is faulty.

Summary of Hab operations: Cleaned and organized hab.

Summary of GreenHab operations: See GreenHab Report.

Summary of ScienceDome operations: None

Summary of RAM operations: Trailer water was cleaned for final time today, remains in the RAM while strong winds persist.

Summary of health and safety issues: None

Questions, concerns, and requests to Mission Support: Doctor Narayanamoorthi recommends biological and chemical testing of water in trailer water tank before clearance for crew use.

Faucet in upper deck of hab was fixed as best we could but it continues to slowly leak.

Sol Summary – April 17th

Crew 192 Sol Summary Report 17Apr2018

Sol 9

Summary Title: Don’t Feed the Penguin

Author: Victoria Varone

Mission Status: Nearing its end, aiming to fit in as much exploration as possible.

Sol Activity Summary: The crew went on two great EVAs today, exploring parts of the Martian terrain that we haven’t yet seen. It’s fascinating to see just how many different kinds of soil, rock, and sediment you can find even in just one area, in varying colors, textures, shapes, and sizes. The winds were strong today (and last night) so extra caution was exercised during EVAs.

Look Ahead Plan: Trying to fit in as many EVAs as possible (pending weather) since weekend weather may not be suitable and our mission ends on Saturday.

Anomalies in Work: None

Weather: Chilly, mix of clouds and sun, strong winds throughout the night and most of the day.

Crew Physical Status: All crew are physically healthy and safe.

EVA: See EVA Report.

Reports to be filed: Sol Summary Report, EVA Report, EVA Request, Photo Report, Commander’s Report, GreenHab Report

Support Requested: None

Maintenance request: See Operations Report

EVA Report – April 17th

EVA Report

EVA #12

17 Apr 2018 – Sol 8

Crew members: Richard Blakeman, Ashok Narayanamoorthi

Objectives:
Conduct an exploratory reconnaissance of 1101 weather permitting. Additionally, continue anti-fogging experiments under varying exertions.

Operations:
0730-1030

Narrative: We proceeded through the route 1101 and were cautiously treading the path in the ATV. We went past the ridges all the way to the fence. Enroute hiked over the ridges and did coms check across the line of sight. Evaluated ATV riding in spacesuit with incremental difficulty in driving.

Vehicles
Blue ATVs #1 and #3

__________

EVA #13

17 April 2018 – Sol 8

Crew Members: Joe Dituri, Andreea Radulescu

Objectives: To explore the east ridge near white moon road 1572 and to the east. There was some indication that part of our science projects crashed there. Also we will be testing two piece vs. one piece suits for longer duration missions.

Operations:
1500-1800

Narrative: We proceeded north on Lowell but did not see a sign for 1572….We missed it but turned around and found what we think was 1572. Explored the east side and this is the most red section similar to what Mars may look like. Great area. We also explored the area slightly south of that road and east. Need to explore north of that road because it gets more dramatic there.

Vehicles
Red ATV and Blue ATV #2

Commander Report – April 17th

Joseph Dituri
17 Apr 18

Title: Just realized I did not file my taxes…ooops (a la Apollo 13) …;*)

Today started with a training and discussion on operational risk management. We use this at every turn before any evolution. We then proceeded to explore south by ATV. Prior to leaving we policed some wind blown scraps from around the HAB and secured them and the final skirt removal pieces on the Mars Society trailer. Upon return the team cleared and cleaned the tank in the RAM with hydrogen peroxide and final wipe with alcohol. When doc gives us go ahead we will return the tank to Mars Society’s trailer and call that cleaning task complete. Likely Wednesday.

Left over peanut butter cookies and dehydrated milk were a big hit upon return along with hot chocolate for the weary EVA goers who worked very hard.

Afternoon EVA explored the northen side of Lowell road up to 1572. There is no sign for road 1572 (route sign is blank) so we missed it. Came back to it and explored east side which is the redest area we have seen to date. We could see some more exciting exploration to the south so we moved south and climbed into east side of Lowell road by Tank wash which was fabulous exploration.

Macaroni and cheese for dinner and we have a great day in total.

IMPRESSED: I was impressed with the doctor’s knowledge and research when it came to decontaminating the water tank. Ashok (Blue 03) was on the spot when it came to rehab of the tank. BZ doc!

LEARNED: We need to schedule science/training in mid day gaps. More of a plan helps the crew pass time since we are short on science. See POA&M

POA&M: Intend to continue exploration of our northern areas tomorrow.

Training scheduled:
1. Psychology and crew selection – Doc and Commander
2. Suborbital mission profile – Rich Blakeman
3. Social Media – Victoria
~ Dituri sends